In a probe that uses the Hall effect to measure magnetic fields, a 12.8 A current passes through a 1.80 cm -wide 1.32 mm -thick strip of sodium metal. If the Hall emf is 1.90 μV , what is the magnitude of the magnetic field (take it perpendicular to the flat face of the strip)? Assume one free electron per atom of Na, and take its specific gravity to be 0.971.
Answer #1

Volume, V = 22.99/0.971 = 23.67 cm^3

Electron density, n = (6.02 x 10^23)/(23.67 x 10^-6) = 2.54 x 10^28 electrons/m^3

Magnetic field, B = (1.90 x 10^-6 x 2.54 x 10^28 x 1.6 x 10^-19 x 0.00132)/12.8

B = 0.80 T

Comment in case any doubt please rate my answer....